OpenStudy (anonymous):

write an equation for each line passing through the given set of points.give final answer in a.slope intercept from & b. standard form. (4,10) (6,12)

OpenStudy (anonymous):

for the eq of line passing through two points use (y-y1)=((y2-y1)/(x2-x1)) *(x-x1)

OpenStudy (anonymous):

thus here it is (y-10)= (12-10)/(6-4) *(x-4) y-10=x-4 x-y=-6 this can be re arranged as y=x+6 (slope intercept form here slope is 1 and y intercept is 6)